1. A policeman on duty detects a drop of 15% in the pitch of the horn of a car as it crosses him. Caalculate the speed of the car, if the velocity of sound is 330 m/s. 


2. Calculate the temperature at which the speed of sound will be 2 times its value at 0 degree celsius.

Grade:Upto college level

1 Answers

vikas askiitian expert
13 years ago

frequency heard by the policemen is 100-15=85% of total

 let n be the freuency of sound emmited by car horn

 we can use ,N=nV/V-v

 here N is appherent frequency ,V velocity of sound ,v velocity of car and n frequency of horn

 v= 58.2m/s

 speed of sound is directly proportional to underoot T

  v=k(t^1/2)

 v1/v2=(t1/t2)^1/2

  T=1092k